# Robust Conformal Outlier Detection under Contaminated Reference Data

### v1

@@ 5 Discussion
However, the experiments also reveal a limitation of our Label-Trim method: while it improves power compared to standard conformal inference, it often remains too conservative, particularly when the labeling budget is very limited, leaving room for further improvement. A promising direction for future research is to enhance Label-Trim with active learning strategies [26, 13, 31], enabling the removal of more outliers without increasing the labeling budget.
A second limitation of the Label-Trim approach is its reliance on actively collecting new annotations. In scenarios where a flexible labeling budget is unavailable but access to a small, clean reference set is feasible, this dependency becomes restrictive. As our experiments demonstrate, the limited sample size imposes a fundamental constraint on the power of conformal outlier detection methods. This raises an intriguing question for future research: given a small clean reference set and a larger, contaminated reference set, how can we effectively and safely clean the contaminated data to enhance detection power at test time?
@@ Acknowledgments
M. S. was partly supported by NSF grant DMS 2210637 and by a Capital One CREDIF Research Award. Y. R. and M. B. were funded by the European Union (ERC, SafetyBounds, 101163414). Views and opinions expressed are however those of the authors only and do not necessarily reflect those of the European Union or the European Research Council Executive Agency. Neither the European Union nor the granting authority can be held responsible for them.
@@ A.2.2 An Alternative View Based on Mixture Distributions
Next, we provide an additional theoretical result concerning the conservativeness of conformal outlier detection methods, to supplement the result presented in Section 2.3 from a point of view closer to that of Sesia et al. [33].
@@
Corollary A.3 reformulates Theorem 1 in Sesia et al. [33] under the setup in (8). This result quantifies the behavior of conformal outlier detection methods in the presence of contaminated data and establishes guarantees on the type-I error rate. This result complements our analysis of the conservativeness of these methods.
@@
The proof adapts the argument of Theorem 1 in Sesia et al. [33] to the outlier detection setting considered here. Specifically, we follow the structure of the original proof, making adjustments to account for the presence of inliers and outliers in the calibration set.
@@ A.3 Validity of the Label-Trim Method
A.3.1 Proof of Theorem 3.1 | Main Steps
@@
Now, consider an imaginary “mirror" version of this method that applies the label-trim algorithm with two key differences:
@@
A.3.2 Proof of Theorem 3.1 | Proof of Equation (9)
@@
  • Suppose $S_{n+1}$ is among the $m+1$ largest scores in $\{S_{i}\}_{i=1}^{n+1}$. This case is illustrated below:
    In this scenario, the mirror trimming approach additionally labels the test score $S_{n+1}$, which, under the null hypothesis, is an inlier and is thus not removed. As a result, both trimming procedures yield the same set; i.e., $\tilde{\operatorname{\mathcal{D}}}_{\mathrm{cal}}^{\mathrm{LT}}=\operatorname{\mathcal{D}}_{\mathrm{cal}}^{\mathrm{LT}}$. Consequently, $\hat{Q}_{1-\alpha}^{\mathrm{LT,n+1}}=\tilde{Q}_{1-\alpha}^{\mathrm{LT,n+1}}$. [Uncaptioned image]
@@
A.3.3 Proof of Theorem 3.1 | Proof of Equation (11)
@@ B.1 Datasets
Dataset Shuttle [3] Credit-card [2] KDDCup99 [1]
@@
For visual datasets, we use the OpenOOD benchmark [43, 39]. Specifically, for each dataset and contamination rate, we perform a one-time training of the ReAct outlier detection model [35], which operates on feature representations extracted from a pre-trained ResNet-18 [43, 19]. The model applies a percentile-based threshold (set to 90%) to truncate activations, where the threshold is computed on the contaminated train set. These truncated activations then pass through the fully connected layer of the pre-trained ResNet-18. The outlier score is computed using an energy-based log-sum-exp function applied to these truncated activations. After training, we save the outlier scores for the remaining outlier samples and the CIFAR-10 test set. We randomly subsample this pool of scores to construct the calibration and test sets, ensuring that all sets are disjoint.
@@
B.2 Tabular Datasets
In this section, we provide additional real-data experiments conducted on the credit card [2] and KDDCup99 datasets, complementing the analysis provided for the shuttle dataset in the main manuscript. Each figure in this section corresponds to and extends the figures presented in the main text.
@@ B.2 Tabular Datasets
In Figure 2 of the main manuscript, we analyze the performance of conformal outlier methods as a function of the contamination rate $r$. Here, we repeat the same experiment on the credit-card and KDDCup99 datasets (Figures 5 and 6). The performance trends are similar to the one presented in the main manuscript: both the Standard and Small-Clean methods achieve valid type-I error rate but exhibit conservative behavior. In contrast, the Naive-Trim method fails to control the type-I error rate. The Label-Trim method attains improved power while practically controlling the type-I error at level $\alpha$.
@@
In Figure 3 of the main manuscript, we evaluate the performance of the Label-Trim, Small-Clean and, Oracle methods as a function of the labeling budget $m$. Figures 7 and 8 extend this analysis to the credit-card and KDDCup99 datasets, respectively. Consistent with the trends observed in Figure 3, increasing the labeling budget improves the performance of the Label-Trim method in terms of both type-I error and power. Notably, although the condition in Theorem 3.1 is no longer satisfied for $m>50$, the Label-Trim method still maintains valid type-I error control at the desired level $\alpha$.
@@
In Figure 4 we examine the performance of conformal outlier detection methods as a function of the target type-I error level $\alpha$. Here, we replicate the experiments on the credit-card and KDDCup99 datasets (Figures 10 and 11). In line with the trends observed in Figure 4, the Label-Trim method performs well at low type-I error rates. Notably, for $\alpha=0.01$, the Label-Trim method outperforms the baselines, while practically controlling the type-I error at level $\alpha$, even though the condition of Theorem 3.1 is not satisfied in this case. This highlights the robustness of our approach.
@@
B.3 Visual Datasets
This section provides detailed results for each dataset, covering various contamination rates and target type-I error levels.
@@ B.3 Visual Datasets
Tables 4(c), 5(c), 6(c), 7(c), 8(c) and 9(c) present the type-I error rates and the power of each method, with the power reported relative to the Standard method (normalized to 1). The actual power of the Standard method is included in the last row of each table.
@@
Table 4: Comparison of conformal outlier detection methods on Texture dataset (outliers) and CIFAR-10 dataset (inliers) for varying contamination rate $r$ and target type-I error level $\alpha$. The empirical power is presented relative to the Standard method (higher is better). Results are averaged across 100 random splits of the data, with standard errors presented in parentheses.
@@
Table 5: Comparison of conformal outlier detection methods on SVHN dataset (outliers) and CIFAR-10 dataset (inliers) for varying contamination rate $r$ and target type-I error level $\alpha$. The empirical power is presented relative to the Standard method (higher is better). Results are averaged across 100 random splits of the data, with standard errors presented in parentheses.
@@
Table 6: Comparison of conformal outlier detection methods on Places365 dataset (outliers) and CIFAR-10 dataset (inliers) for varying contamination rate $r$ and target type-I error level $\alpha$. The empirical power is presented relative to the Standard method (higher is better). Results are averaged across 100 random splits of the data, with standard errors presented in parentheses.
@@
Table 7: Comparison of conformal outlier detection methods on MNIST dataset (outliers) and CIFAR-10 dataset (inliers) for varying contamination rate $r$ and target type-I error level $\alpha$. The empirical power is presented relative to the Standard method (higher is better). Results are averaged across 100 random splits of the data, with standard errors presented in parentheses.
@@
Table 8: Comparison of conformal outlier detection methods on CIFAR-100 dataset (outliers) and CIFAR-10 dataset (inliers) for varying contamination rate $r$ and target type-I error level $\alpha$. The empirical power is presented relative to the Standard method (higher is better). Results are averaged across 100 random splits of the data, with standard errors presented in parentheses.
@@
Table 9: Comparison of conformal outlier detection methods on TinyImageNet dataset (outliers) and CIFAR-10 dataset (inliers) for varying contamination rate $r$ and target type-I error level $\alpha$. The empirical power is presented relative to the Standard method (higher is better). Results are averaged across 100 random splits of the data, with standard errors presented in parentheses.

### v2

@@ 5 Discussion
However, the experiments also reveal a limitation of our Label-Trim method: while it improves power compared to standard conformal inference, it often remains too conservative, particularly when the labeling budget is very limited, leaving room for further improvement. A promising direction for future research is to enhance Label-Trim with active learning strategies (Makili et al., 2012; Fannjiang et al., 2022; Prinster et al., ), enabling the removal of more outliers without increasing the labeling budget.
While we provide robustness results for calibration under contaminated reference data, our analysis assumes that the inlier calibration and test points are drawn i.i.d. from the same distribution. An important direction for future work is to extend this setting beyond the i.i.d. assumption, possibly by building on ideas from Tibshirani et al. (2019); Podkopaev & Ramdas (2021); Sesia et al. (2024); Barber et al. (2023).
Another limitation of the Label-Trim approach is its reliance on actively collecting new annotations. In scenarios where a flexible labeling budget is unavailable but access to a small, clean reference set is feasible, this dependency becomes restrictive. As our experiments demonstrate, the limited sample size imposes a fundamental constraint on the power of conformal outlier detection methods. This raises an intriguing question for future research: given a small clean reference set and a larger, contaminated reference set, how can we effectively and safely clean the contaminated data to enhance detection power at test time?
@@
Another direction related to the above discussion is how to account for uncertainty in the labeling process. In our current formulation, we assume that all calibration points are labeled as inliers, but some may in fact be outliers. Importantly, we have no indication of which points are mislabeled, nor any signal of uncertainty in the labels. Exploring how to incorporate such uncertainty into our framework could enhance its practical utility. The line of work presented in Stutz et al. (2023); Javanmardi et al. (2024); Caprio et al. (2025) may offer a valuable starting point for such an extension: it introduces techniques for handling ambiguous labels, though in the context of multi-class classification rather than outlier detection.
Notably, our paper also provides practical guidance on how to annotate data in situations where there is uncertainty about whether a point is an inlier or an outlier. As indicated by Lemma 2.2 and supported by our empirical results, when there is uncertainty about a point’s label, treating it as an inlier is a conservative strategy that preserves type-I error control. This observation suggests a compelling connection between label ambiguity (Stutz et al., 2023; Javanmardi et al., 2024; Caprio et al., 2025) and contamination in the reference set, which merits further investigation.
@@ Acknowledgments
M. S. was partly supported by NSF grant DMS 2210637 and by a Capital One CREDIF Research Award. Y. R. and M. B. were funded by the European Union (ERC, SafetyBounds, 101163414). Views and opinions expressed are however those of the authors only and do not necessarily reflect those of the European Union or the European Research Council Executive Agency (ERCEA). Neither the European Union nor the granting authority can be held responsible for them. This research was also partially supported by the Israel Science Foundation (ISF grant 729/21). Y. R. acknowledges additional support from the Career Advancement Fellowship at the Technion.
@@ A.2.2 An Alternative View Based on Mixture Distributions
Next, we provide an additional theoretical result concerning the conservativeness of conformal outlier detection methods, to supplement the result presented in Section 2.3 from a point of view closer to that of Sesia et al. (2024).
@@
Corollary A.3 reformulates Theorem 1 in Sesia et al. (2024) under the setup in (8). This result quantifies the behavior of conformal outlier detection methods in the presence of contaminated data and establishes guarantees on the type-I error rate. This result complements our analysis of the conservativeness of these methods.
@@
The proof adapts the argument of Theorem 1 in Sesia et al. (2024) to the outlier detection setting considered here. Specifically, we follow the structure of the original proof, making adjustments to account for the presence of inliers and outliers in the calibration set.
@@ A.3 Validity of the Label-Trim Method
A.3.1 Proof of Theorem 3.1 — Main Steps
@@
Now, consider an imaginary “mirror” version of this method that applies the label-trim algorithm with two key differences:
@@
A.3.2 Proof of Theorem 3.1 — Proof of Equation (9)
@@
  • Suppose $S_{n+1}$ is among the $m+1$ largest scores in $\{S_{i}\}_{i=1}^{n+1}$. This case is illustrated below: In this scenario, the mirror trimming approach additionally labels the test score $S_{n+1}$, which, under the null hypothesis, is an inlier and is thus not removed. As a result, both trimming procedures yield the same set; i.e., $\tilde{\operatorname{\mathcal{D}}}_{\mathrm{cal}}^{\mathrm{LT}}=\operatorname{\mathcal{D}}_{\mathrm{cal}}^{\mathrm{LT}}$. Consequently, $\hat{Q}_{1-\alpha}^{\mathrm{LT,n+1}}=\tilde{Q}_{1-\alpha}^{\mathrm{LT,n+1}}$. [Uncaptioned image]
@@
A.3.3 Proof of Theorem 3.1 — Proof of Equation (11)
@@ B.1 Datasets
Dataset Shuttle (shu, ) Credit-card (cre, ) KDDCup99 (KDD, )
@@
For visual datasets, we use the OpenOOD benchmark (Zhang et al., 2024; Yang et al., 2022). Specifically, for each dataset and contamination rate, we perform a one-time training of an outlier detection model, which operates on feature representations extracted from a pre-trained backbone. We include results with the following model configurations:
  • ReAct with ResNet-18: ReAct (Sun et al., 2021) operates on feature representations extracted from a pre-trained ResNet-18 model (Zhang et al., 2024; He et al., 2016). The model applies a percentile-based threshold (set to 90%) to truncate activations, where the threshold is computed on the contaminated train set. These truncated activations then pass through the fully connected layer of the pre-trained ResNet-18. The outlier score is computed using an energy-based log-sum-exp function applied to these truncated activations.
  • ReAct with VGG-19: Same as above, but with a pre-trained VGG-19 backbone (Chen, ) instead of ResNet-18.
  • SCALE with ResNet-18: SCALE (Xu et al., 2024) operates on feature representations extracted from a pre-trained ResNet-18 model (Zhang et al., 2024; He et al., 2016). The model rescales the activations using a sample-specific factor, defined as the sum of all activations divided by the sum of activations below a certain percentile (set to 65%). Similar to ReAct, the outlier score is computed using an energy-based log-sum-exp function applied to the rescaled activations.
After training, we save the outlier scores for the remaining outlier samples and the CIFAR-10 test set. We randomly subsample this pool of scores to construct the calibration and test sets, ensuring that all sets are disjoint.
@@
B.2 Additional Experiments
B.2.1 Tabular Datasets
In this section, we provide additional real-data experiments conducted on the credit card (cre, ) and KDDCup99 datasets, complementing the analysis provided for the shuttle dataset in the main manuscript. Figures 6, 7, 8, 9, 10, 11 and 12 corresponds to and extends the figures presented in the main text.
We further report supplementary results across all three tabular datasets considered in this work, including: additional outlier detection models (Figures 13 and 14); higher contamination levels (Figure 15); strategic outlier injection (Figures 16, 17, 18 and 19); and test-time distributional shifts in the outlier population (Figure 20).
@@ B.2.1 Tabular Datasets
In Figure 3 of the main manuscript, we analyze the performance of conformal outlier methods as a function of the contamination rate $r$. Here, we repeat the same experiment on the credit-card and KDDCup99 datasets (Figures 6 and 7). The performance trends are similar to the one presented in the main manuscript: both the Standard and Small-Clean methods achieve valid type-I error rate but exhibit conservative behavior. In contrast, the Naive-Trim method fails to control the type-I error rate. The Label-Trim method attains improved power while practically controlling the type-I error at level $\alpha$.
@@
In Figure 4 of the main manuscript, we evaluate the performance of the Label-Trim, Small-Clean and, Oracle methods as a function of the labeling budget $m$. Figures 8 and 9 extend this analysis to the credit-card and KDDCup99 datasets, respectively. Consistent with the trends observed in Figure 4, increasing the labeling budget improves the performance of the Label-Trim method in terms of both type-I error and power. Notably, although the condition in Theorem 3.1 is no longer satisfied for $m>50$, the Label-Trim method still maintains valid type-I error control at the desired level $\alpha$.
@@
In Figure 5 we examine the performance of conformal outlier detection methods as a function of the target type-I error level $\alpha$. Here, we replicate the experiments on the credit-card and KDDCup99 datasets (Figures 11 and 12). In line with the trends observed in Figure 5, the Label-Trim method performs well at low type-I error rates. Notably, for $\alpha=0.01$, the Label-Trim method outperforms the baselines, while practically controlling the type-I error at level $\alpha$, even though the condition of Theorem 3.1 is not satisfied in this case. This highlights the robustness of our approach.
@@
Results with additional outlier detection models
Here, we consider two additional outlier detection models: Local Outlier Factor (LOF) with 100 estimators and One-Class Support Vector Machine (OC-SVM) with an RBF kernel, both implemented via scikit-learn. Figures 13 and 14 present the performance of all methods when applied with LOF and OC-SVM, respectively. As can be seen, the proposed Label-Trim controls the type-I error across both models. When the outlier detection model is less discriminative (e.g., LOF or OC-SVM on the KDDCup99 dataset), the candidate set becomes less effective and power decreases. Nevertheless, our method improves over the baselines, including the Small-Clean method.
(a) shuttle
(b) credit-card
(c) KDDCup99
Figure 13: Comparison of conformal outlier detection methods on real datasets as a function of the contamination rate $r$. All methods utilize a Local Outlier Factor model. Other details are as in Figure 3 in the main manuscript.
(a) shuttle
(b) credit-card
(c) KDDCup99
Figure 14: Comparison of conformal outlier detection methods on real datasets as a function of the contamination rate $r$. All methods utilize a One-Class SVM model. Other details are as in Figure 3 in the main manuscript.
Results with higher contamination rate
Figure 15 shows the performance of conformal outlier detection methods as a function of the contamination rate $r$, evaluated at higher levels than those considered in the main manuscript. The performance trend aligns with that observed at lower contamination rates, across all datasets. Both the Standard and Small-Clean methods control the type-I error at the nominal level but remain overly conservative. In contrast, Naive-Trim fails to control the type-I error. The proposed Label-Trim method achieves improved power while controlling the type-I error at level $\alpha$.
(a) shuttle
(b) credit-card
(c) KDDCup99
Figure 15: Comparison of conformal outlier detection methods on real tabular datasets as a function of the contamination rate $r$. Other details are as in Figure 3 in the main manuscript.
Results with strategic outlier injection
We extend the experimental setting from the main manuscript by evaluating different outlier injection strategies. Instead of injecting outliers at random, we selected outliers that more closely resemble inliers—specifically, outliers whose nonconformity scores fall below a given score percentile. Figure 16 illustrates the nonconformity scores for both outliers and inliers under different outlier injection strategies for the “shuttle” dataset, highlighting that lower-percentile outliers increasingly resemble inliers. This sets the stage to evaluate the performance of our proposed method under these challenging settings. Figure 17 presents the performance of all conformal methods on the same dataset and outlier injection strategies, as a function of the target type-I error level. Following that figure, we see that Label-Trim controls the type-I error while improving power. Similar trends hold on the credit-card and KDDCup99 datasets (Figures 18 and 19).
(a) Random outlier injection
(b) $70$th percentile
(c) $50$th percentile
Figure 16: Histogram of nonconformity scores for inliers and outliers in a contaminated calibration subset of the “shuttle” data, with a contamination rate of 5% for different outlier injection strategies.
(a) Random outlier injection
(b) $70$th percentile
(c) $50$th percentile
Figure 17: Performance on a real dataset “shuttle” as a function of the target type-I error rate $\alpha$ for different outlier injection strategies. Other details are as in Figure 5 in the main manuscript.
(a) Random outlier injection
(b) $70$th percentile
(c) $50$th percentile
Figure 18: Performance on a real dataset “credit-card” as a function of the target type-I error rate $\alpha$ for different outlier injection strategies. Other details are as in Figure 5 in the main manuscript.
(a) Random outlier injection
(b) $90$th percentile
(c) $80$th percentile
(d) $70$th percentile
Figure 19: Performance on a real dataset “KDDCup99” as a function of the target type-I error rate $\alpha$ for different outlier injection strategies. Other details are as in Figure 5 in the main manuscript.
Results with test-time drifting outliers
We further investigate the robustness of the proposed method under test-time distribution shift in the outlier population. Using the injection strategies described earlier, we simulate a shift on the “shuttle” dataset where the outlier distribution gradually changes over time. Specifically, while the training and calibration sets contain high-percentile outliers, the test set progressively includes more challenging, low-percentile outliers. As shown in Figure 20, Label-Trim maintains type-I error control throughout the distribution shift, outperforming all baselines and approaching oracle performance.
Figure 20: Performance on a real dataset “shuttle” as a function of drift in the outlier test distribution. Outliers in the training and calibration sets are drawn from the $30$th to $100$th percentile of the outlier score distribution, while the test set varies across different percentile ranges. Other details are as in Figure 3 in the main manuscript.
B.2.2 Visual Datasets
@@
We also report performance averaged across the six datasets for two additional models: ReAct (Sun et al., 2021) with a VGG-19 backbone and SCALE (Xu et al., 2024) with a ResNet-18 backbone, shown in Tables 4(c) and 5(c), respectively. These results align with our main findings and further demonstrate the effectiveness of the proposed Label-Trim method.
Following this, we provide detailed results for each dataset, reporting type-I error rates and the power of each method, with the power reported relative to the Standard method (normalized to 1). The actual power of the Standard method is included in the last row of each table. Specifically, Tables 6(c), 7(c), 8(c), 9(c), 10(c) and 11(c) present the detailed per-dataset results corresponding to Tables 1(b) and 3(a). Additionally, detailed per-dataset results for the additional models are provided in Tables 12(c), 13(c), 14(c), 15(c), 16(c) and 17(c) and Tables 18(c), 19(c), 20(c), 21(c), 22(c) and 23(c), corresponding to Table 4(c) and Table 5(c), respectively.
Table 4: Comparison of conformal outlier detection methods on six visual datasets for varying contamination rate $r$ and target type-I error level $\alpha$. All methods utilize the ReAct (Sun et al., 2021) method with a pretrained VGG-19. Other details are as in Table 1(b) in the main manuscript.
Contamination rate
1% 3% 5%
Method Power Type-I Error Power Type-I Error Power Type-I Error
Standard 1.0 ($\pm$ 0.0317) 0.009 ($\pm$ 0.0004) 1.0 ($\pm$ 0.0377) 0.006 ($\pm$ 0.0003) 1.0 ($\pm$ 0.0422) 0.004 ($\pm$ 0.0002)
Oracle (infeasible) 1.095 ($\pm$ 0.033) 0.01 ($\pm$ 0.0004) 1.368 ($\pm$ 0.0418) 0.01 ($\pm$ 0.0004) 1.633 ($\pm$ 0.0488) 0.009 ($\pm$ 0.0003)
Naive-Trim (invalid) 1.464 ($\pm$ 0.0373) 0.018 ($\pm$ 0.0005) 2.5 ($\pm$ 0.0531) 0.031 ($\pm$ 0.0006) 3.447 ($\pm$ 0.0592) 0.04 ($\pm$ 0.0008)
Label-Trim 1.095 ($\pm$ 0.033) 0.01 ($\pm$ 0.0004) 1.364 ($\pm$ 0.0418) 0.01 ($\pm$ 0.0004) 1.606 ($\pm$ 0.0487) 0.009 ($\pm$ 0.0003)
(a) Target type-I error rate $\alpha=0.01$
Contamination rate
1% 3% 5%
Method Power Type-I Error Power Type-I Error Power Type-I Error
Standard 1.0 ($\pm$ 0.0251) 0.018 ($\pm$ 0.0005) 1.0 ($\pm$ 0.027) 0.014 ($\pm$ 0.0004) 1.0 ($\pm$ 0.0291) 0.01 ($\pm$ 0.0004)
Oracle (infeasible) 1.068 ($\pm$ 0.0247) 0.021 ($\pm$ 0.0005) 1.221 ($\pm$ 0.0297) 0.02 ($\pm$ 0.0005) 1.37 ($\pm$ 0.0323) 0.019 ($\pm$ 0.0006)
Naive-Trim (invalid) 1.252 ($\pm$ 0.0256) 0.028 ($\pm$ 0.0006) 1.759 ($\pm$ 0.0326) 0.039 ($\pm$ 0.0007) 2.212 ($\pm$ 0.0345) 0.048 ($\pm$ 0.0008)
Small-Clean 0.813 ($\pm$ 0.0593) 0.017 ($\pm$ 0.0019) 0.593 ($\pm$ 0.0737) 0.011 ($\pm$ 0.0018) 0.352 ($\pm$ 0.0663) 0.005 ($\pm$ 0.0013)
Label-Trim 1.056 ($\pm$ 0.0244) 0.02 ($\pm$ 0.0005) 1.177 ($\pm$ 0.0292) 0.019 ($\pm$ 0.0005) 1.253 ($\pm$ 0.0312) 0.016 ($\pm$ 0.0005)
(b) Target type-I error rate $\alpha=0.02$
Contamination rate
1% 3% 5%
Method Power Type-I Error Power Type-I Error Power Type-I Error
Standard 1.0 ($\pm$ 0.0202) 0.028 ($\pm$ 0.0006) 1.0 ($\pm$ 0.0238) 0.022 ($\pm$ 0.0005) 1.0 ($\pm$ 0.0247) 0.017 ($\pm$ 0.0005)
Oracle (infeasible) 1.074 ($\pm$ 0.02) 0.031 ($\pm$ 0.0006) 1.194 ($\pm$ 0.0256) 0.03 ($\pm$ 0.0007) 1.318 ($\pm$ 0.0271) 0.029 ($\pm$ 0.0007)
Naive-Trim (invalid) 1.186 ($\pm$ 0.0202) 0.037 ($\pm$ 0.0007) 1.51 ($\pm$ 0.0248) 0.048 ($\pm$ 0.0008) 1.848 ($\pm$ 0.0272) 0.056 ($\pm$ 0.0009)
Small-Clean 0.709 ($\pm$ 0.045) 0.019 ($\pm$ 0.0018) 0.85 ($\pm$ 0.0541) 0.021 ($\pm$ 0.002) 0.932 ($\pm$ 0.0568) 0.019 ($\pm$ 0.002)
Label-Trim 1.048 ($\pm$ 0.0202) 0.03 ($\pm$ 0.0006) 1.118 ($\pm$ 0.0248) 0.027 ($\pm$ 0.0006) 1.167 ($\pm$ 0.026) 0.023 ($\pm$ 0.0006)
(c) Target type-I error rate $\alpha=0.03$
Table 5: Comparison of conformal outlier detection methods on six visual datasets for varying contamination rate $r$ and target type-I error level $\alpha$. All methods utilize the SCALE (Xu et al., 2024) method with a pretrained ResNet-18. Other details are as in Table 1(b) in the main manuscript.
Contamination rate
1% 3% 5%
Method Power Type-I Error Power Type-I Error Power Type-I Error
Standard 1.0 ($\pm$ 0.0358) 0.009 ($\pm$ 0.0003) 1.0 ($\pm$ 0.0375) 0.006 ($\pm$ 0.0003) 1.0 ($\pm$ 0.0407) 0.004 ($\pm$ 0.0002)
Oracle (infeasible) 1.144 ($\pm$ 0.037) 0.01 ($\pm$ 0.0003) 1.472 ($\pm$ 0.0474) 0.01 ($\pm$ 0.0003) 1.708 ($\pm$ 0.0522) 0.01 ($\pm$ 0.0004)
Naive-Trim (invalid) 1.736 ($\pm$ 0.0414) 0.017 ($\pm$ 0.0004) 2.93 ($\pm$ 0.0553) 0.028 ($\pm$ 0.0006) 4.144 ($\pm$ 0.0614) 0.038 ($\pm$ 0.0007)
Label-Trim 1.144 ($\pm$ 0.0371) 0.01 ($\pm$ 0.0003) 1.456 ($\pm$ 0.0469) 0.01 ($\pm$ 0.0003) 1.627 ($\pm$ 0.0501) 0.009 ($\pm$ 0.0003)
(a) Target type-I error rate $\alpha=0.01$
Contamination rate
1% 3% 5%
Method Power Type-I Error Power Type-I Error Power Type-I Error
Standard 1.0 ($\pm$ 0.0236) 0.018 ($\pm$ 0.0004) 1.0 ($\pm$ 0.0267) 0.013 ($\pm$ 0.0004) 1.0 ($\pm$ 0.0288) 0.01 ($\pm$ 0.0003)
Oracle (infeasible) 1.076 ($\pm$ 0.0244) 0.021 ($\pm$ 0.0005) 1.306 ($\pm$ 0.0279) 0.02 ($\pm$ 0.0005) 1.583 ($\pm$ 0.0316) 0.02 ($\pm$ 0.0005)
Naive-Trim (invalid) 1.247 ($\pm$ 0.0253) 0.027 ($\pm$ 0.0006) 1.86 ($\pm$ 0.0294) 0.036 ($\pm$ 0.0006) 2.532 ($\pm$ 0.0335) 0.047 ($\pm$ 0.0008)
Small-Clean 0.831 ($\pm$ 0.0598) 0.018 ($\pm$ 0.0019) 0.609 ($\pm$ 0.0765) 0.011 ($\pm$ 0.0017) 0.348 ($\pm$ 0.0744) 0.005 ($\pm$ 0.0014)
Label-Trim 1.062 ($\pm$ 0.0243) 0.02 ($\pm$ 0.0005) 1.226 ($\pm$ 0.0284) 0.018 ($\pm$ 0.0005) 1.368 ($\pm$ 0.0315) 0.015 ($\pm$ 0.0004)
(b) Target type-I error rate $\alpha=0.02$
Contamination rate
1% 3% 5%
Method Power Type-I Error Power Type-I Error Power Type-I Error
Standard 1.0 ($\pm$ 0.0201) 0.027 ($\pm$ 0.0006) 1.0 ($\pm$ 0.0212) 0.021 ($\pm$ 0.0005) 1.0 ($\pm$ 0.022) 0.016 ($\pm$ 0.0004)
Oracle (infeasible) 1.082 ($\pm$ 0.0204) 0.031 ($\pm$ 0.0006) 1.233 ($\pm$ 0.0227) 0.029 ($\pm$ 0.0006) 1.398 ($\pm$ 0.0237) 0.03 ($\pm$ 0.0006)
Naive-Trim (invalid) 1.19 ($\pm$ 0.0197) 0.036 ($\pm$ 0.0006) 1.547 ($\pm$ 0.0229) 0.044 ($\pm$ 0.0007) 1.9 ($\pm$ 0.0239) 0.055 ($\pm$ 0.0009)
Small-Clean 0.712 ($\pm$ 0.045) 0.019 ($\pm$ 0.0019) 0.862 ($\pm$ 0.055) 0.02 ($\pm$ 0.002) 0.971 ($\pm$ 0.0606) 0.021 ($\pm$ 0.0021)
Label-Trim 1.05 ($\pm$ 0.0204) 0.029 ($\pm$ 0.0006) 1.116 ($\pm$ 0.022) 0.025 ($\pm$ 0.0006) 1.178 ($\pm$ 0.0224) 0.022 ($\pm$ 0.0005)
(c) Target type-I error rate $\alpha=0.03$
@@
Table 6: Comparison of conformal outlier detection methods on Texture dataset (outliers) and CIFAR-10 dataset (inliers) for varying contamination rate $r$ and target type-I error level $\alpha$. All methods utilize the ReAct (Sun et al., 2021) method with a pretrained ResNet-18. The empirical power is presented relative to the Standard method (higher is better). Results are averaged across 100 random splits of the data, with standard errors presented in parentheses.
@@
Table 7: Comparison of conformal outlier detection methods on SVHN dataset (outliers) and CIFAR-10 dataset (inliers) for varying contamination rate $r$ and target type-I error level $\alpha$. All methods utilize the ReAct (Sun et al., 2021) method with a pretrained ResNet-18. The empirical power is presented relative to the Standard method (higher is better). Results are averaged across 100 random splits of the data, with standard errors presented in parentheses.
@@
Table 8: Comparison of conformal outlier detection methods on Places365 dataset (outliers) and CIFAR-10 dataset (inliers) for varying contamination rate $r$ and target type-I error level $\alpha$. All methods utilize the ReAct (Sun et al., 2021) method with a pretrained ResNet-18. The empirical power is presented relative to the Standard method (higher is better). Results are averaged across 100 random splits of the data, with standard errors presented in parentheses.
@@
Table 9: Comparison of conformal outlier detection methods on MNIST dataset (outliers) and CIFAR-10 dataset (inliers) for varying contamination rate $r$ and target type-I error level $\alpha$. All methods utilize the ReAct (Sun et al., 2021) method with a pretrained ResNet-18. The empirical power is presented relative to the Standard method (higher is better). Results are averaged across 100 random splits of the data, with standard errors presented in parentheses.
@@
Table 10: Comparison of conformal outlier detection methods on CIFAR-100 dataset (outliers) and CIFAR-10 dataset (inliers) for varying contamination rate $r$ and target type-I error level $\alpha$. All methods utilize the ReAct (Sun et al., 2021) method with a pretrained ResNet-18. The empirical power is presented relative to the Standard method (higher is better). Results are averaged across 100 random splits of the data, with standard errors presented in parentheses.
@@
Table 11: Comparison of conformal outlier detection methods on TinyImageNet dataset (outliers) and CIFAR-10 dataset (inliers) for varying contamination rate $r$ and target type-I error level $\alpha$. All methods utilize the ReAct (Sun et al., 2021) method with a pretrained ResNet-18. The empirical power is presented relative to the Standard method (higher is better). Results are averaged across 100 random splits of the data, with standard errors presented in parentheses.